15. Write the verbal sentence as an equation. Then solve the eqaution
Twelve less than 2 times a number is equal to 15 minus 7 times the number
Can someone help with this equation?

Answer by longjonsilver(2297)

let the number be a, so:
Twelve less than 2 times a number is equal to 15 minus 7 times the number
12 less than 2a = 15 - 7a

so, 12 less than 2a:
what is 12 less than 20? answer is found by 20-12
what is 12 less than 30? answer is found by 30-12
so, what is 12 less than 2a? answer is found by 2a-12

so, 2a-12 = 15-7a is the equation.

2a-12 = 15-7a --> add 7a to both sides
9a-12 = 15 --> add 12 to both sides
9a = 27
a = 27/9
a = 3

cheers
jon